OpenStudy (anonymous):

the measues of two angles are 16z-9 and 4z+3. find the measures of the angles.

OpenStudy (anonymous):

could be anything. there must be something missing

OpenStudy (anonymous):

yea maybe a diagram or something!

OpenStudy (anonymous):

of which shape or diagram

OpenStudy (anonymous):

Are they complementary, supplementary, or vertical?

OpenStudy (anonymous):

yea its complementary

OpenStudy (anonymous):

if i'm not wrong somwhere then z=4.8

OpenStudy (anonymous):

z=1

OpenStudy (anonymous):

cause im trying to find out what Z is

OpenStudy (radar):

Add them up, make the sum = 90 degrees, then solve for z.

OpenStudy (radar):

16z-9+4z+3=90 20z-6=90 20z=96 z=96/20=4.8 just like Ishaan94 stated.

OpenStudy (anonymous):

lol.. i dunno why i posted z= 1.. i made them equal :O my bad :P